# Random Exclusion Codes: Quantum Advantages of Single-Shot Communication

### v2

@@ Abstract
Identifying information tasks that offer advantages over their classical counterparts by exploiting quantum resources finds the usefulness of quantum information technologies. In this work, we introduce a two-party communication primitive, random exclusion code (REC), which is a single-shot prepare-and-measure protocol where a sender encodes a random message into a shorter sequence and a receiver attempts to exclude a randomly chosen letter in the original message. We present quantum advantages in RECs in two ways: probability and dimension. We show that RECs with quantum resources achieve higher success probabilities than classical strategies. We verify that the quantum resources required to describe detection events of RECs have a smaller dimension than classical ones. We also show that a guessing counterpart, random access codes (RACs), may not have a dimension advantage over classical resources. Our results elucidate various possibilities of achieving quantum advantages in two-party communication.
@@
One of the key challenges in Quantum Information Theory is to address information tasks that demonstrate the distinct nature of quantum and classical resources. A gap appearing from the distinction, even if it is small, may be amplified toward practical applications, namely, quantum advantages in communication Winter (2013). It is hence of both fundamental and practical importance to identify a simplest information scenario, such as a point-to-point communication protocol, such that quantum advantages are present. The Holevo theorem is the first attempt along this line, considering a channel capacity with i.i.d. quantum states, but showing no quantum advantage Holevo (2011, 1998); Schumacher and Westmoreland (1997).
@@
It is random access codes (RACs) that have verified the existence of quantum advantages in two-party communication Ambainis et al. (2009, 2024); Tavakoli et al. (2015); Carmeli et al. (2020). We emphasize that it is a single-shot scenario in which a single outcome is obtained from a detection event, from which a receiver learns about a message. Then, the guess is correct with some probability. The quantum advantage is demonstrated by the fact that the guessing probability in RACs is higher with quantum resources than classical strategies.
@@
Although the connection is not direct, RACs are closely related to quantum state discrimination Helstrom (1969), see also Bae and Kwek (2015); Barnett and Croke (2009); Bergou (2007). Interestingly, quantum state anti-discrimination, which efficiently excludes a state from an ensemble, has been used as an effective tool to clarify the distinctions between quantum and classical resources in a single-shot experiment Caves et al. (2002); Pusey et al. (2012); Bandyopadhyay et al. (2014a); Heinosaari and Kerppo (2018). Quantum advantages can be obtained in terms of a success probability Uola et al. (2020); Ducuara et al. (2020), and a quantum-classical gap is also present in detection events. For instance, the Pusey-Barrett-Rudolph (PBR) theorem Pusey et al. (2012) rules out an impossible interpretation of quantum states by deriving a detection event that cannot happen with quantum resources.
@@
Upon examining the fundamental tasks closely, it becomes clear that discrimination and exclusion are distinct in nature. Perfect discrimination of states achieves perfect exclusion by assigning an item from perfect guessing to another one. The converse does not generally hold. Thus, quantum state exclusion is operationally more general. Then, state exclusion may lead to even greater possibilities of quantum advantages.
@@
As discussed above, efforts to distinguish quantum and classical resources in a single-shot scenario take one of the approaches. One is comparing the success probability for a single detection event; the probability with quantum resources is higher than that of classical ones, e.g., RACs. The other is to take into account the detection events. For instance, the PBR theorem signifies a detection event that is impossible to occur with quantum resources but is possible with independent ontic preparations Pusey et al. (2012); Harrigan and Spekkens (2010). Note that the Bell theorem is comparable to the latter in that it also rules out an ontic interpretation of quantum states; in contrast to the PBR result, two parties repeat a Bell experiment for estimating joint probabilities Bell (1964); Brunner et al. (2014). Then, it is worth mentioning that the quantum-classical gap from the Bell theorem has been developed as a practical tool for device-independent quantum information processing, such as randomness extraction Pironio et al. (2010), quantum key distribution Acín et al. (2007), state certification Šupić and Bowles (2020); Šupić et al. (2021), etc.
@@
We show quantum advantages of RECs in two ways: one is the probability advantage for a single detection event, and the other is the dimension advantage for a collection of possible detection events. Firstly, a higher success probability in the exclusion task of a REC can be achieved with quantum resources than with classical ones. Secondly, quantum resources in a smaller Hilbert space achieve RECs that can be realized only in a larger classical dimension. Note that a classical dimension denotes the minimal number of distinguishable letters of an alphabet.
@@
The importance of the results is twofold. On the one hand, we relate state exclusion, a fundamental tool for verifying distinctions between the elements of quantum and classical theories Caves et al. (2002); Bandyopadhyay et al. (2014b); Crickmore (2020); Pusey et al. (2012), to a communication primitive that may find practical applications. On the other hand, we demonstrate the quantum-classical gap within a single-shot framework by highlighting the differences between quantum and classical supports for detection events, see also Pusey et al. (2012). We also find that a RAC may not have a dimension advantage.
@@
We begin with notations for an $(n,m)$ REC. A message of Alice is written by $a=a_{1}...a_{n}$ where each is chosen from a set of $m$ distinguishable letters, $a_{i}\in\{0,\cdots,m-1\}$. An encoding prepares a $d$-dimensional system, and let $b_{j}$ denote Bob’s decoding of $a_{j}$ once he decides to read out $a_{j}$, see also Fig. 1. With an REC, Bob wants to find $b_{j}$ that is not equal to $a_{j}$. In this case, the probability of successfully excluding a correct value on average is given,
@@
In contrast to RECs, Bob for RACs wants to guess Alice’s message. The average success probability is found,
In both cases above in Eqs. (1) and (2), parties aim to maximize the success probability in all strategies with given resources.
Note that state discrimination and state exclusion are equivalent for an alphabet of two letters: guessing one immediately excludes the other, and vice versa. For example, for binary messages $a_{i}\in\{0,1\}$, Bob is with two-state discrimination and it holds that
@@
where the maximization runs over all strategies $S$. From above, one immediately concludes that RECs bear quantum advantages of RACs in the $(2,2)$ instance. However, the relation above no longer holds when an alphabet contains more than two letters; excluding a correct value does not yield a single outcome, but rather multiple ones.
@@
Let us then consider the next simplest case, the $(2,3)$ REC, where two-trit messages $a=a_{1}a_{2}$ for $a_{i}\in\{0,1,2\}$ are encoded to a binary system, for which a quantum realization is a single qubit. We then demonstrate that Bob has a higher probability of excluding a message sent by Alice with quantum resources than classical ones.
@@
We first compute the average success probability with classical resources. The maximum is obtained as follows,
$\max_{C}p^{(\mathrm{REC})}=\frac{8}{9},$ (4)
where $C$ denotes all classical strategies. A detailed analysis is provided in the Appendix. An optimal classical strategy is to encode messages $a\in\{00,\cdots,22\}$ based on the classification,
$\mathbb{P}_{0}=\{11,12,21,22\}\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ \mathbb{P}_{1}=\{00,01,02,10,20\}.$
For $a\in\mathbb{P}_{0}$, the encoding of Alice is $c=0$, which is sent to Bob. For $a\in\mathbb{P}_{1}$, it is $c=1$. Then, Bob applies his decoding $b_{i}=c$ for $i=1,2$. This classical strategy achieves an average success probability above.
We next present a $(2,3)$ REC with single qubit states in the following,
where
$|\psi_{02}\rangle$ $=$ $\frac{|0\rangle+|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ |\psi_{12}\rangle=\frac{|0\rangle-|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ \rho_{22}=\mathds{1}/2.$
We construct a decoding with two-outcome measurements,
$\mathcal{M}_{1}=\{M_{10},M_{11},M_{12}\}\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ \mathcal{M}_{2}=\{M_{20},M_{21},M_{22}\}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ (6)
where
and |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$ and |±$|\pm i\rangle=(|0\rangle\pm i|1\rangle)/\sqrt{2}$. Notice that the states and measurements above are on a half-plane in a Bloch sphere.
The quantum protocol works as follows. Alice sends a state $\rho_{a_{1}a_{2}}$ and Bob first decides which one, $a_{1}$ or $a_{2}$, he is going to read. For the first trit $a_{1}$, he applies measurement $\mathcal{M}_{1}$, and for the second, $a_{2}$, he applies $\mathcal{M}_{2}$. The probability of successfully excluding a one-bit message of Alice is given by,
$p^{\mathrm{(REC)}}_{Q}$ $=$ $\frac{1}{18}\sum_{a}\sum_{i=1}^{2}{\rm Tr}[(M_{i(a_{i}\oplus 1)}+M_{i(a_{i}\oplus 2)})\leavevmode\nobreak\ \rho_{a_{1}a_{2}}]\leavevmode\nobreak\ \leavevmode\nobreak$ (7)
where $\oplus$ represents addition modulo three. It follows that
$p^{\mathrm{(REC)}}_{Q}=\frac{7+\sqrt{2}}{9}$ (8)
which is larger than the classical value in Eq. (4). Thus, the usefulness of quantum resources in RECs is shown.
There may exist states and measurements other than those in Eqs. (5) and (6), in particular non-zero positive-operator-value-measure (POVM) elements, such that the success probability is even higher. That is, one can ask if the bound in Eq. (8) is maximal. Our answer is in the affirmative. Firstly, we find that to maximize the success probability with qubit transmission, it suffices to restrict states and measurements to a half-plane on a Bloch sphere. This also holds true even if parties aim at a guessing task, and hence, applies to RACs with qubits. Secondly, extensive numerical optimization of the average success probability finds Eq. (8) with two POVM elements are zeros. We detail the derivations in the Appendix. The gap between quantum and classical bounds is,
which is about $0.046$.
For comparisons, let us also consider a $(2,3)$ RAC with single-qubit transmission. An upper bound to success probabilities with classical strategies is known as $5/9$ Saha et al. (2023). We compute the maximum average probability with quantum resources, applying the aforementioned result of reducing to a half-plane. We find that a relabeling of our quantum protocol for the REC gives the quantum maximum for the RAC; messages $a=a_{1}a_{2}$ are encoded in the states orthogonal to ones given in Eq. (5) and the same measurements as in Eq. (6) are applied for decoding. The quantum bound is obtained as $(4+\sqrt{2})/9$. Interestingly, the gap is identical to Eq. (9), i.e.,
$\Delta_{QC}^{(\mathrm{RAC})}=\Delta_{QC}^{(\mathrm{REC})}$ (10)
for the considered $(2,3)$ RAC and $(2,3)$ REC.
In what follows, we are motivated by the observation that a single-shot scenario gives an outcome among possible events. We investigate the minimal quantum and classical resources required to describe possible detection events in the $(2,3)$ REC, and present a dimensional advantage of quantum resources over classical ones. To this end, we exploit communication matrices, a tool that characterizes detection events of a two-party protocol.
A communication matrix is defined for an input $\alpha$ and an outcome $\beta$ as follows Heinosaari and Kerppo (2019); Frenkel and Weiner (2015); Heinosaari et al. (2020),
$C_{\alpha\beta}={\rm Prob}_{M|P}(\beta|\alpha)$
where $P$ and $M$ denote preparation and measurement, respectively. For instance, a communication matrix of a guessing task for an alphabet of two letters in a noiseless manner corresponds to a diagonal matrix, $S_{2}=\mathrm{diag}[1,1]$. As for an exclusion task for three letters $\{a,b,c\}$, the communication matrix with the unit success probability is given as
$A_{3}=\frac{1}{2}\begin{bmatrix}0&1&1\\ 1&0&1\\ 1&1&0\end{bmatrix},$ (11)
meaning that once a letter is sent, say $a$, a receiver decodes others, i.e., $b$ or $c$, with equal probabilities $1/2$. In the following, we investigate the quantum and classical dimensions required for constructing a communication matrix. We recall that a quantum dimension is referred to as the Hilbert space, and a classical one is defined by the number of distinguishable letters.
Figure 2: Trine states (black) in Eq. (12) and an anti-trine measurement (blue) in Eq. (13) can realize an unbiased exclusion task, see its communication matrix in Eq. (11). For instance, an outcome $\pi_{0}$ concludes state $\psi_{1}$ or state $\psi_{2}$ with equal probabilities $1/2$ excluding $\psi_{0}$ unambiguously.
To illustrate the approach for quantifying minimal dimensions, let us consider a communication matrix in Eq. (11). A quantum realization can be constructed with trine states, see also Fig. 2
and anti-trine measurements,
$\pi_{\beta}=\frac{2}{3}|{\psi}_{\beta}^{\perp}\rangle\langle{\psi}_{\beta}^{\perp}|\,\,\,{\beta}=0,1,2$ (13)
where $\langle\psi_{k}|{\psi}_{\beta}^{\perp}\rangle=0$. States and measurements above can reproduce a communication matrix in Eq. (11); the Hilbert space of quantum resources has dimension two. Classically, however, it is necessary to have three distinguishable letters to describe the matrix in Eq. (11) as the matrix is of rank three. Therefore, a dimensional advantage of quantum resources is shown.
In general, a communication matrix may have different cardinalities for inputs and outputs. Let us consider an $n\times m$ communication matrix $C$. The minimal dimension of classical resources corresponds to the nonnegative rank, denoted by $\mathrm{rank}_{+}(C)$, as a minimal value of $k$ such that there exist an $n\times k$ nonnegative matrix $L$ and a $k\times m$ nonnegative matrix $R$ that satisfy $LR=C$ Cohen and Rothblum (1993). As for the quantum case, it has been shown that the minimal dimension of quantum resources is given by the positive semidefinite (psd) rank, denoted by $k=\mathrm{rank}_{psd}({C})$, defined as the smallest integer $k$ such that there exists a positive semidefinite $k\times k$ matrices $A_{1},\ldots,A_{n}$ and $B_{1},\ldots,B_{m}$ satisfying $C_{ij}={\rm Tr}[A_{i}B_{j}]$ Fawzi et al. (2015). In the following, we compute the nonnegative and semidefinite ranks of a communication matrix.
The communication matrix for the considered $(2,3)$ REC with the unit success probability, such that exclusion is realized in an unbiased manner, can be obtained as follows,
$D_{3}$ $=\frac{1}{4}\begin{bmatrix}0&0&0&0&1&1&0&1&1\\ 0&0&0&1&0&1&1&0&1\\ 0&0&0&1&1&0&1&1&0\\ 0&1&1&0&0&0&0&1&1\\ 1&0&1&0&0&0&1&0&1\\ 1&1&0&0&0&0&1&1&0\\ 0&1&1&0&1&1&0&0&0\\ 1&0&1&1&0&1&0&0&0\\ 1&1&0&1&1&0&0&0&0\\ \end{bmatrix}.$ (14)
It turns out that we have $D_{3}=A_{3}\otimes A_{3}$ and it follows that we have $\mathrm{rank}_{+}(D_{3})=9$ Heinosaari et al. (2020).
A minimal quantum dimension can be found by upper and lower bounds to the positive-semidefinite rank. We consider a quantum realization to ensure an upper bound and then compute a general lower bound. We show that they coincide.
A quantum realization using two qubits is as follows. Alice encodes $a_{1}$ of her string into one qubit, and $a_{2}$ into another, as one of the three trine states in Eq. (12). On receiving the systems, Bob applies an anti-trine measurement (see Eq. (13)) onto the relevant qubit and can hence exclude the state unambiguously. Furthermore, due to the symmetry in the ensemble, the remaining outcomes occur equiprobably. A two-qubit space with a dimension of four, therefore, suffices for the purpose. We have
$\mathrm{rank}_{psd}(D_{3})\leq 4.$ (15)
In Ref. Lee et al. (2017), a lower bound to psd ranks for positive doubly stochastic matrices is shown. Applying the bound to $D_{3}$, we have
$\mathrm{rank}_{psd}({D_{3}})\geq\max_{q}\frac{1}{\sum_{i,j}q_{i}q_{j}F(\mathrm{col}_{i}(D_{3}),\mathrm{col}_{j}(D_{3}))^{2}}$
where $\mathrm{col}_{i}(D_{3})$ is the $i$-th column of $D_{3}$, the maximum runs over probability vectors $q=[q_{1},\cdots,q_{9}]$ and $F(a,b)=\sum_{i}\sqrt{a_{i}b_{i}}$. We compute that the lower bound is
$\mathrm{rank}_{psd}({D_{3}})\geq 4,$
which occurs if $q_{i}=1/9$ for $i=1,\cdots,9$. As upper and lower bounds coincide, we conclude a minimal quantum dimension $4$. The $(2,3)$ REC can be realized on a smaller quantum dimension than a classical one, showing a distinction in quantum and classical supports for detection events.
We also consider quantum and classical dimensions for a $(2,3)$ RAC protocol. Notice that the matrix for noiseless communication is given by
$\mathbf{1}_{9}=\mathrm{diag}[1,1,1,1,1,1,1,1,1]$
which is on the support of a dimension $9$. For this matrix, it can be shown that $\mathrm{rank}_{+}(\mathbf{1}_{9})=\mathrm{rank}_{psd}(\mathbf{1}_{9})=9$ Fawzi et al. (2015). Hence, no dimensional advantage is achieved between the quantum and classical dimensions in the RAC. In other words, if a protocol can implement random access perfectly, we can implement exclusion perfectly by relabeling outcomes. The converse, however, is not true: one cannot realize random access with certainty on the quantum support where random exclusion is perfect.
Finally, RECs can be generalized to higher dimensions. We detail in the Appendix the analysis about the $(2,m)$ REC with a single qubit and show quantum advantages for all values of $m\geq 2$. Namely, we have the maximum of the success probability with classical resources,
$\max_{C}p^{(REC)}=1-\frac{1}{m^{2}}.$
We extend the quantum strategy considered for $m=3$ to a larger $m$ and have,
$p_{Q}^{(REC)}=1-\frac{2-\sqrt{2}}{m^{2}},$
which is strictly higher than the classical one. Thus, quantum advantages remain for all $m\geq 2$.
In conclusion, we have presented a framework for identifying quantum advantages of a single-shot protocol by introducing a communication primitive, RECs, which relies on an exclusion task. A single-shot scenario has a detection event where outcomes occur probabilistically. Quantum advantages are provided by a probability and a dimension. Firstly, the probability that the exclusion task is successful from a single detection event is higher with quantum strategies than with classical ones. Secondly, by exploiting communication matrices, we have shown that the quantum resources of a smaller dimension can realize RECs with a larger alphabet. The dimensional quantum advantage may be lacking in RACs; a protocol for a RAC implies an implementation of a REC by relabeling, but the converse does not hold, similarly to the distinction between state discrimination and state exclusion.
In future directions, it would be interesting to investigate general properties of $(n,m)$ RECs and their relations to quantum state exclusion. In particular, we leave it open to characterize optimal measurements for RECs and quantum state exclusion to find general relations between them. We also conjecture that RECs may have quantum advantages for all $n,m\geq 2$. In addition, it is also interesting to combine the probability and the dimension advantages of quantum resources to amplify the usefulness. We envisage that our results would lead to practical quantum information applications that outperform their classical counterparts.
This work is supported by Business Finland funded project BEQAH, the National Research Foundation of Korea (Grant No. NRF-2021R1A2C2006309, NRF-2022M1A3C2069728, RS-2024-00408613, RS-2023-00257994) and the Institute for Information & Communication Technology Promotion (IITP) (RS-2023-00229524, RS-2025-02304540).
Appendix A Optimal success with classical resource in $(2,3)$ REC
@@
The maximum classical average success probability in the simplest random exclusion task, consisting of length $n=2$ words made from size $m=3$ alphabet and dimension of the communicated system $d=2$, is $8/9$.

### v3

@@ Abstract
Useful applications of quantum information technologies can be found by identifying tasks in which quantum resources outperform their classical counterparts. In this work, we introduce a two-party communication primitive, random exclusion code (REC), which is a single-shot prepare-and-measure protocol where a sender encodes a random message into a shorter sequence and a receiver attempts to exclude a randomly chosen letter in the original message. We present quantum advantages in RECs in two ways: probability and dimension. We show that RECs with quantum resources achieve higher success probabilities than classical strategies. We verify that the quantum resources required to describe detection events of RECs have a smaller dimension than classical ones. We also show that a guessing counterpart, random access codes (RACs), may not have a dimension advantage over classical resources. Our results elucidate various possibilities of achieving quantum advantages in two-party communication.
I Introduction
@@
One of the key challenges in quantum information theory is to address information tasks that demonstrate the distinct nature of quantum and classical resources. A gap in performance appearing from the distinction, even if it is small, may be amplified toward practical applications, namely, quantum advantages in communication winter. It is hence of both fundamental and practical importance to identify a simplest information scenario, such as a point-to-point communication protocol, such that quantum advantages are present. The Holevo theorem, concerning the channel capacity with i.i.d. quantum states, was the first attempt along these lines but failed to show a quantum advantage Holevo2011; 651037; PhysRevA.56.131.
@@
It is random access codes (RACs) that have verified the existence of quantum advantages in two-party communication Ambainis2009; Ambainis2024RAC; Tavakoli2015; carmeli2020quantum. We emphasize that the setting is a single-shot scenario, in which a single outcome is obtained from a detection event, from which a receiver learns about a message. The guess is then correct with some probability. The quantum advantage is demonstrated by the fact that the guessing probability in RACs is higher if quantum, rather than classical, resources are used.
@@
Although the connection is not direct, RACs are closely related to quantum state discrimination Helstrom:1969aa; Bae_2015; Barnett:09; Bergou:2007aa. Interestingly, quantum state exclusion, also known as anti-discrimination, which efficiently excludes a state from an ensemble, has been used as an effective tool to clarify the distinctions between quantum and classical resources in single-shot experiments Caves2002; PuseyPBR2012; PhysRevA.89.022336; HeKe18. Quantum advantages can be obtained in terms of a success probability PhysRevLett.125.110402; PhysRevResearch.2.033374, and a quantum-classical gap is also present in detection events. The Pusey-Barrett-Rudolph (PBR) theorem PuseyPBR2012, for instance, rules out a certain interpretation of quantum states by pinpointing a detection event that cannot happen with quantum resources.
@@
Upon examining the fundamental tasks closely, it becomes clear that discrimination and exclusion are distinct in nature. Perfect discrimination of states may achieve perfect exclusion by reassigning events to different labels. The converse does not generally hold. Quantum state exclusion is thus operationally more general. State exclusion may therefore lead to greater possibilities of quantum advantages.
@@
As discussed above, efforts to distinguish quantum and classical resources in a single-shot scenario take one of several approaches. One is to compare the success probability for a single detection event. As with RACs, the probability with quantum resources may be higher than that with classical resources. The other approach is to take into account detection events. Along these lines, the PBR theorem signifies a detection event that is impossible with quantum resources but is possible with independent ontic preparations PuseyPBR2012; Harrigan:2010aa. Note that the Bell theorem is comparable to the latter in that it also rules out an ontic interpretation of quantum states; in contrast to the PBR result, two parties repeat a Bell experiment for estimating joint probabilities PhysicsPhysiqueFizika.1.195; Brunner2014. It is worth mentioning that the quantum-classical gap from the Bell theorem has been developed as a practical tool for device-independent quantum information processing, such as randomness extraction Pironio:2010aa, quantum key distribution PhysRevLett.98.230501 and state certification Supic2020selftestingof; Supic2021deviceindependent. One would therefore expect useful applications to follow from other quantum-classical distinctions.
@@
We show the quantum advantages of RECs in two ways: one is the probabilistic advantage for a single detection event, and the other is the dimension advantage for a collection of possible detection events. Firstly, a higher success probability in the exclusion task of a REC can be achieved with quantum resources than with classical ones. Secondly, quantum resources in a smaller Hilbert space achieve RECs that can be realized only in a larger classical dimension. Note that a classical dimension denotes the minimal number of distinguishable letters of an alphabet.
@@
The importance of the results is twofold. On the one hand, we relate state exclusion, a fundamental tool for verifying distinctions between the elements of quantum and classical theories Caves2002; Bandyopadyhyay2014Exclusion; Crickmore2020Elimination; PuseyPBR2012, to a communication primitive that may find practical applications. On the other hand, we demonstrate the quantum-classical gap within a single-shot framework by highlighting the differences between quantum and classical supports for detection events, see also PuseyPBR2012. We also find that RACs may not have a comparable dimension advantage.
The article is structured as follows. We first remind the reader of some relevant results in the theory of quantum state exclusion and random access coding. We then introduce random exclusion codes and show that quantum resources outperform classical ones in the simplest non-trivial setting. We then analyse the required dimensional supports for perfect quantum and classical RECs and see that the former require fewer dimensions. Throughout, our results are contrasted with those from the equivalent RAC.
II Background
II.1 Quantum state exclusion
In information processes, one is typically interested in determing that a datum takes a particular value from a known, discrete alphabet. Exclusion conversely refers to the opposite task of determining that a datum does not have a given value Crickmore2020Elimination. The latter task is in general easier than the former, but recent works have shown that perfect implementation of this task is a signature of quantum behaviour.
In quantum state exclusion, a classical alphabet of size $m$ will be associated with a set of quantum states $\{\rho_{\alpha}\}_{\alpha=0}^{m-1}$ by a party Alice. She is given a value $\alpha$ at random and sends a system prepared in the appropriate state to a receiving party, Bob, who must announce one of the values which $\alpha$ is not. He determines his choice by implementing an $m$-outcome probability operator valued measurement (POVM) $\{\pi_{\beta}\}_{\beta=0}^{m-1}$.
We can distinguish between different criteria for state exclusion. Perfect state exclusion is that with zero-error, so that, whenever Alice prepares the state $\alpha$, Bob is guaranteed to announce a value $\beta\neq\alpha$. Uniform state discrimination, on the other hand, refers to the case in which all allowed outcomes occur with uniform probability. We will return to this point in later sections of the manuscript.
Figure 1: Trine states (black) in Eq. (1) and an anti-trine measurement (blue) in Eq. (2) can realize an unbiased exclusion task. For instance, an outcome $\pi_{0}$ concludes state $\psi_{1}$ or state $\psi_{2}$ with equal probabilities $1/2$ while excluding $\psi_{0}$ unambiguously.
The simplest non-trivial construction which is able to implement perfect and uniform state exclusion is the set of trine states
and anti-trine measurements
$\pi_{\beta}=\frac{2}{3}|{\psi}_{\beta}^{\perp}\rangle\langle{\psi}_{\beta}^{\perp}|,\,\,\,{\beta}=0,1,2$ (2)
where $\langle\psi_{\beta}|{\psi}_{\beta}^{\perp}\rangle=0$. This set of states and measurements is shown in Fig. 1 and has been used to implement key distribution Phoenix2000 and in showing gaps between classical and quantum behaviour Heinosaari2024Simple. The defining feature in this example is symmetry. In similar symmetric constellations one finds the excluding POVM as a group covariant POVM HeKe18.
The problem of state exclusion was first studied by Caves et al. Caves2002 and came to wider attention when it was used to prove the PBR theorem regarding the $\psi$-onticity of quantum theory. Subsequent investigations have aimed to characterise the set of quantum ensembles which states can be unambiguously excluded Johnston2025tightbounds, and related conclusive exclusion to various foundational problems in quantum theory Havlicek2020Exclusion; Srikumar2024contextuality; Stratton2024Choi and the construction of zero-error communication channels Duan2016; Chiribella2025. The key takeaway from these results has been that perfect state exclusion is a signature of quantum behaviour and, furthermore, may be exploited for improved communications.
II.2 Random Access Codes
Random access codes are a communication primitive widely studied in the context of quantum theory for both their information-theoretic applications and foundational significance Ambainis2009; Ambainis2024RAC; Tavakoli2015; carmeli2020quantum. Alice holds a string $a=a_{1}...a_{n}$ consisting of $n$ independent classical labels each taken equiprobably from an $m$-symbol alphabet $\{0,\cdots,m-1\}$. This data is encoded into a system, either classical or quantum, of dimension $d$ which is then sent to the receiver, Bob. Bob’s task is to read out some subset of the information. He is given a question label, $i$, and aims to respond with $b_{i}$ such that $b_{i}=a_{i}$. Measures of his ability to perform this task use either the worst-case probability of success or the average success probability, given by
If the communication system is quantum, Alice’s encoding will map the data to a set of density matrices $\{\rho_{a}\}$ and Bob’s decoding will be a set of positive operator-valued measures (POVMs) $M_{i}$. If the system is instead classical, then the encoding and decoding are simply maps to and from dits. In this article, we are particularly interested in the case in which $n=2$ and the quantum and classical systems to be compared are both two-dimensional. In this restricted setting, upper bounds are known to be Tavakoli2015; Ambainis2024RAC:
$\begin{split}\max_{Q}p^{(RAC)}&=\frac{1}{2}(1+\frac{1}{\sqrt{m}})\\ \max_{C}p^{(RAC)}&=\frac{1}{2}(1+\frac{1}{m}).\end{split}$ (4)
where the former optimisation is over quantum strategies and the latter over classical ones. In this case, and in general, it holds that $\max_{Q}p^{(RAC)}>\max_{C}p^{(RAC)}$.
III Random Exclusion Codes
@@
We begin with the notation for an $(n,m)$ REC. A message of Alice’s is written by $a=a_{1}...a_{n}$ where each is chosen from a set of $m$ distinguishable letters, $a_{i}\in\{0,\cdots,m-1\}$. An encoding prepares a $d$-dimensional system, and we let $b_{j}$ denote Bob’s decoding of $a_{j}$ once he decides to read out $a_{j}$, see also Fig. 2. With an REC, Bob wants to find $b_{j}$ that is not equal to $a_{j}$. In this case, the probability of successfully excluding a correct value on average is given,
@@
For RACs, in contrast to RECs, Bob wants to guess Alice’s message and the average success probability is computed from Eq. (3). In both the tasks, above in Eqs. (3) and (5), parties aim to maximize the success probability over all strategies with given resources.
Note that state discrimination and state exclusion are equivalent for an alphabet of two letters: guessing one immediately excludes the other, and vice versa. For example, for binary messages $a_{i}\in\{0,1\}$, Bob can apply two-state discrimination and it holds that
@@
where the maximization runs over all strategies $S$. From this, one immediately concludes that RECs bear quantum advantages of RACs in the $(2,2)$ instance. However, the relation above no longer holds when an alphabet contains more than two letters as excluding a correct value does not yield a single outcome but rather multiple ones.
@@
Let us then consider the next simplest case, the $(2,3)$ REC, where two-trit messages $a=a_{1}a_{2}$ for $a_{i}\in\{0,1,2\}$ are encoded to a binary system, for which a quantum (classical) realization is a single qubit (bit). In what follows, we demonstrate a quantum advantage by showing that Bob has a higher probability of excluding a message sent by Alice with quantum rather than classical resources.
IV Quantum advantage in $(2,3)$-RECs
We prove analytically that, if the communication system is a classical bit, the maximum value of the average probability of success in the considered task is $\frac{8}{9}$. We then show that if instead quantum systems of dimension two are used, a larger value of $\frac{7+\sqrt{2}}{9}$ can be attained. It is worth mentioning here that for the considered REC, since the alphabet $\{0,1,2\}$ is non-binary, the corresponding RAC task is not equivalent, which can be seen from the fact that the maximum classical value for the RAC task is $\frac{5}{9}$ (see Result 3 in the Ref. deba+PRA2023). The inequivalence also holds for the quantum values, where the maximum success probability for the corresponding RAC is $\frac{4+\sqrt{2}}{9}$.
IV.1 Optimal bit implementation
@@
We first compute the maximum average success probability with classical resources, which can be summarised as follows:
@@
The maximum classical average success probability in the $(2,3)$ random exclusion task, consisting of length $n=2$ words using a size $m=3$ alphabet and $d=2$ dimension communicated system, is $8/9$.